OpenStudy (anonymous):

Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form.

OpenStudy (anonymous):

well you have to know the slope formula and point slope formula

OpenStudy (ybarrap):

\( (y2 - y1) / (x2 - x1) = m\) , plug in your coordinates and see what ya get

OpenStudy (anonymous):

first you have to get the slope using the formula @ybarrap put

OpenStudy (anonymous):



OpenStudy (anonymous):

this answer should give you your slope = m you can use this to plug it into the next equation which will be the point slope formula y - y1 =m ( x - x1 )

OpenStudy (anonymous):

just in case you were wondering though the i used y -y1 instead of y1 - y2 its only a way distinguish between the coordinates

OpenStudy (anonymous):

so... just plug in you coordinates again make sure you dont get mixed up in you calculations and that should give the answer in slope intercept form